OpenStudy (anonymous):

(x-8)(3x+7)=0 solve the equation

jimthompson5910 (jim_thompson5910):

Use the idea that if AB = 0, then either A = 0 or B = 0 So in the case of (x-8)(3x+7)=0 either x-8 = 0 OR 3x+7=0 Solve each equation for x to find your answers